wiki:UC20180205

Version 1 (modified by trac, 6 months ago) ( diff )

--

APO 3.5-m Users Committee Phonecon, 2/05/18

Attending: Nancy Chanover (NMSU), Mark Klaene (APO), Rachel Kuzio de Naray (GSU), Kal Kadlec (UW), Sarah Tuttle (UW), Mary Beth Kaiser (JHU), Bill Ketzeback (APO), John Bally (CU), Ben Williams (UW), Anne Verbiscer (UVa; minute taker), Russet McMillan (APO), John Tobin (OU), Leslie Hebb (NAPG), Jon Holtzman (NMSU), Chip Kobulnicky (UWy)


User feedback and comments from institutional representatives:

  • Georgia State - nothing to report
  • JHU - two issues: (a) lost time on recent run due to echelle software - had to be rebooted and rerun (this is discussed in more detail below), (b) if their members are not logged into a JHU machine they are unable to access/read the Users Committee minutes on the wiki (Nancy will discuss this with Jon Holtzman, and until it gets resolved she will also include the minutes in an email the way it used to be handled)
  • UVA - nothing to report
  • NAPG - nothing to report
  • Oklahoma - some of their members just had a good night with TripleSpec, but 2 weeks ago a user had telescope tracking problem while using Agile. It was fixed but then the data were not easy to use. Telescope was not guiding so exposures were taken with just telescope tracking. (this is discussed in more detail below)
  • NMSU - nothing to report
  • Colorado - nothing to report
  • Wyoming - no report
  • Seattle - no report
  • Washington - no report

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 12/06/17 through 1/30/18

0) Overview

Reasonably good weather this period, a couple small snow storms but a lot of clear dry nights. DIS and Echelle scattered light being monitored closely; both holding at less than desirable levels although Echelle has now risen up to our lower threshold. DIS vacuum is being serviced this week. We had planned a earlier DIS vacuum service run but NIC-FPS ICC problems delayed that.

1) Telescope

Telescope was operational with minimal issues. We did see some minor communications issues that we are monitoring and looking into some hardware issues.

2) Instruments

DIS: We have pumped the red camera and blue camera is on the pump in an effort to clear any contamination. DCam gave us shutter issues despite it being a extremely high duty cycle rated shutter. Also we are still getting some banded images. Original DCam is at Andor for servicing.

TripleSpec has been operational.

Agile has been operational this period with one minor issue.

Echelle is operational. Inter-Order Light is worse than we like but still improving slightly. Ecam shutter (new spare installed a couple weeks ago) is causing some issues. Original shutter is out for refurbishment. Estimate return date is TBD from manufacturer. There is also readout banding that is causing some issues. A replacement camera order from Fingerlakes has been ordered.

NIC-FPS is operational but the old ICC hardware died and was out of service for a few weeks. Fortunately we had other old hardware that we were able to make a useable ICC from.

ARCTIC is operational . We did have a unplanned warmup due to loss of cryotiger gas. The system has been recharged but we will need time once we get to warm weather to try and locate a leak.

3) ARCSAT - SurveyCam was returned from Andor. It does have a bad row in the center that we can not do much with. The telescope is now down for major maintenance. This includes repolishing and coating M1 and M2, Dome and telescope PM , looking at M2 focusing mechanism, software improvements, and possibly new guider.


Additional telescope and instrument discussion:

DIS: We pumped on both the red and blue cameras. Red camera got 3 days back of servicing. For the blue camera, we tried a short pump to remove contamination; that removed some of it. Longer pumping has removed around 50%. Russet will add better numbers. By the numbers it looked like it improved the scattered light down to about a fourth of where it was before. Still assessing what is the rate of increase since red was finished earlier. Rate of increase on red was much slower than what it was in the autumn. On the blue side the pumping was not as successful as when we did it in November. As long as rate of increase is low, we're better off than we were. Expect it to get bad again when the weather warms up; it seems to be worse this year than in previous years. Russet has done a good job of constructing the schedule so DIS can be taken offline occasionally to work to reduce the contamination issue. We're not going to be kicking people off the instrument to take it down. Russet reminded everyone that scheduling is coming up, and asked that institutional schedulers keep an eye out for DIS requests during bright time; we will have trouble supporting that since we typically do the DIS pumping during bright time.

Agile: Regarding the tracking issue, Russet said it was a simple problem and simple fix, but it has been rare so not everyone remembers the fix. Russet will remind the obs-specs of the fix on the next engineering telecon and document it.

Echelle: there was a number of updates contained in Mark's report. The readout banding issue is something that at this point we're monitoring. Won't have anything new to report on this until we get the new Finger Lakes camera in for ecam. Mary Beth's readout issue appears to be something different: it cost a significant amount of time for 2 nights in a row. It has not recurred since then and we will continue to monitor it to see how common this problem is. The nature of the problem is that the saddlebag electronics stopped talking to the Echelle ICC program so it wasn't talking to TUI and able to take images. Doing the reset is fairly complicated but we're ready to do this should the problem occur again.

Mark commented that several of the above discussion items highlight the fact that our instrumentation suite is aging and we need to worry about their long term support. We will continue to monitor this issue at the observatory level.

ARCSAT: The two mirrors (M1 and M2) did go out for repolishing. Hopefully they arrived in Georgia and have been hand-carried home to where they will be worked on (this hand-carry transport saved APO a few thousand $!). The telescope will be offline for a couple months; the observatory staff plans to make progress on other ARCSAT-related engineering tasks while the telescope is down.


Q1 ARCSAT scheduling:

Not much to say; telescope was in use for only 1 week in Q1 before it went down.


Q1 3.5m scheduling:

We just published an updated Q1 schedule yesterday. We have an additional B half night in Feb. that we're eager to give away. A halves in March are going quickly.


Q2 3.5-m scheduling:

Ben sent out institutional allocations last week. Requests are due on Feb. 27.


Q2 ARCSAT scheduling:

There has been no call issued for Q2 proposals yet. Since we're not sure when the telescope will be back together, we will hold off on issuing a call until we have an update. When the mirrors come back the telescope will need to be collimated and we've never done that, so it will require a fair amount of time.


TDAWG Update

The Time Domain Astronomy Working Group (TDAWG) is continuing to function well. They have generated a lot of input for the final report, which is pretty close to complete. Ben will send a Doodle poll out and get most of members online to go over it one more time; then their initial task will be done. He hopes to have it finished up this month. Users Committee reps should coordinate with their institutional schedulers regarding requested TDAWG input (concerning the degree to which they would be willing to be preempted for time domain followup work, and/or whether they would be interested in pooling time across ARC institutions to dedicate to time domain followup observations). Ben is awaiting input from two institutions, and he encourages everyone to send in whatever input they have so that it can incorporated into the report.


NA2 Baffle Project

John Wilson submitted a proposal for the NA2 Baffle Project to Nancy, and it has been given a green light to proceed. Previous calls got vignetting information from Joanne that was needed. Joanne provided those files, and Nancy put them on DropBox and put a link to them on the wiki so they are available to anyone who wants to access them. Nothing more to report at this time.


Optical Spectrograph Plan Document

Nancy sent around a document last week summarizing the optical spectrograph strategy. It describes a plan for how to best meet the needs of the ARC user community. The new spectrograph is still in the works, but it has become clear that developing a single new spectrograph that meets all needs of our diverse community will be challenging if not impossible. We have an exciting opportunity with KOSMOS from Kitt Peak. It is being taken off the Mayall 4m telescope to make room for DESI, and NOAO will let us use KOSMOS and modify it for use on the 3.5m for no cost. Nancy would like to move forward with asking the ARC Board of Governors to approve this plan. Comments? Mary Beth reported that feedback from the JHU users has been neutral to positive. Tobin: all positive. One user used it before. Would be good for slitviewer to be added. Anne adds input from John Wilson email expressing support for bringing KOSMOS to APO. Desirements state that we can switch gratings in. Will limit bandpasses available. John Bally asked about high resolution. Mostly interested in specific emission lines. Looks like there are 3 spaces where gratings could be fit in. Is there capability to tilt the grating to tune the wavelength? Sarah thinks this is possible. Optical design requires grism to do this. KOSMOS has a switchable optical design. Not a big deal to do this. Would like to forward this on to ARC Board of Governors and get an MOU signed between NOAO and ARC and get the instrument put in Sarah's lab. Will continue to seek input from the users to see what the impact of bringing this instrument to APO will be. Will proceed from here.


ACTION ITEMS:

open from previous meetings:

  • Russet will iterate with the UVa team to clarify some of the language in the TripleSpec User's Guide. Status: CLOSED. Russet was working with Mike Skrutskie and John Wilson to solve this. A typo has been fixed, but Russet said typo wasn't related to the confusion. Mike thinks the language is sufficiently clear. Do we need feedback from other users? Chip was the original person involved. The issue is the identification of when a pixel is saturated -- with power sampling it can be difficult to tell. Russet determined that this does not need to be an action item with this committee, and will continue to address this offline.
  • Ben and Nancy will discuss the ARCTIC flats issue on an upcoming 3.5m engineering telecon to figure out what is acceptable for the Observing Specialists, and come up with a policy regarding how late the observers can keep them to take flats. Status: CLOSED. The upshot is that we want the obs-specs and users to communicate early and often concerning calibration needs. This specific request is now included in the reminder emails that go out to observers three days before their runs. Users should notify obs-specs what their calibration needs are. Everyone agreed that the reminder emails to observers are very valuable and good feedback comes from those. There was some concern that adding in more detail (e.g. about calibration needs) might be counterproductive since observers may not read all of the content. It was suggested that these emails be made as concise as possible, bulletized and prioritized, with changes highlighted at the top.
  • For the NA2 baffle project, Joanne will send a flat field image around to work with. Based on the information from Joanne Nancy will work with John Wilson to add the amplitude of the current loss of signal in the corners of the ARCTIC FOV to the NA2 Baffle Project Description document. Status: CLOSED. All of the relevant files were transferred to Nancy, and she posted a link to them on the ARCTIC users wiki page.
  • All UC reps: coordinate with your institutional scheduler to get a sense from your users concerning the degree to which they would be willing to be preempted for time domain followup work, and/or whether they would be interested in pooling time across ARC institutions to dedicate to time domain followup observations. Status: OPEN. Ben is still awaiting input from two institutions.

new from this meeting:

None.


Next meeting:

The meeting schedule for the spring semester will be the first Monday of the month at 11:00 am Mountain Time. The next call will thus be March 5, 2018.

Future meetings are scheduled for:

  • April 2
  • May 7
  • June 4
  • July 9

We will evaluate everyone's availability (as well as the telecon schedule) once we get into summer months since we will need to contend with summer travel schedules as well as summer shutdown activities.

Note: See TracWiki for help on using the wiki.